OpenStudy (anonymous):

A hot air balloon holds 3,249 cubic meters of helium. The density of helium is 0.1785 kilograms per cubic meter. How many kilograms of helium does the balloon contain, rounded to the nearest tenth of a kilogram? 1,873.4 kg 2,194.6 kg 2,721.2 kg 579.9 kg

OpenStudy (anonymous):

Use\[D = \frac{ m }{ V }\]Can you rearrange this formula for mass, m?

OpenStudy (anonymous):

I think you would have to multiply them together and you'll get your answer

OpenStudy (anonymous):

Mass= d*v?

OpenStudy (anonymous):

Correct. Just put in the values for density (D) and volume (V) and solve. What do you get?

OpenStudy (anonymous):

d*v= 579.9465

OpenStudy (anonymous):

Great job. Round off to the nearest tenth of a kg and you've got your answer.

OpenStudy (anonymous):

579.9 so D?

OpenStudy (anonymous):

Well done.